OpenStudy (anonymous):

I am going to fail plz help me i am begging i will fan and metal if you help Two lines, A and B, are represented by the equations given below: Line A: y = x − 2 Line B: y = 3x + 4 Which of the following shows the solution to the system of equations and explains why? (−3, −5), because the point does not lie on any axis (−3, −5), because the point satisfies both equations (−4, −8), because the point satisfies one of the equations (−4, −8), because the point lies between the two axes

OpenStudy (anonymous):

plz help me!

Vocaloid (vocaloid):

Didn't I already answer this for you? Line A y = x − 2 Line B: y = 3x + 4 x - 2 = 3x + 4 solve for x, then substitute that value for x to get y

OpenStudy (anonymous):

i told you i didnt get it!

Vocaloid (vocaloid):

what are you still confused about?

OpenStudy (anonymous):

i just dont get it

Vocaloid (vocaloid):

do you know how to solve x - 2 = 3x + 4?

OpenStudy (anonymous):

i just want to know how to solve it and have it explained it to me

OpenStudy (z4k4r1y4):

 what does that tell you about the two lines?

Vocaloid (vocaloid):

well, let's break it down step by step: x - 2 = 3x + 4 what happens when you add 2 to each side?

OpenStudy (anonymous):

i dont know what you mean Do you mean: 2x-2=3x=6?

Vocaloid (vocaloid):

no, that's multiplication ADD 2 to each side of the equal sign (+2)

Vocaloid (vocaloid):

x - 2 + 2 = 3x + 4 + 2 see how I added + 2 to each side?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

great! x - 2 + 2 = x, do you agree?

OpenStudy (anonymous):

umm idk it doesnt look right because what did you do with the 3

Vocaloid (vocaloid):

I am only looking at the left side of the equal sign x - 2 + 2 = 3x + 4 + 2 JUST look at the left side of the equals sign: x - 2 + 2 do you agree that x - 2 + 2 = x?

OpenStudy (anonymous):

yes i do

Vocaloid (vocaloid):

good! so now we have: x = 3x + 4 + 2, do you agree? we replace the left side with x because x - 2 + 2 = x

OpenStudy (anonymous):

yes okay i am starting to get it a little bit

Vocaloid (vocaloid):

ok, now just look at the right side: 3x + 4 + 2 do agree that 3x + 4 + 2 = 3x + 6?

OpenStudy (anonymous):

yes i do

Vocaloid (vocaloid):

awesome, now we put both sides together: x = 3x + 6 but we're not done yet

Vocaloid (vocaloid):

what happens when you subtract 3x from each side?

OpenStudy (anonymous):

would it become x=3x+3x+6?

Vocaloid (vocaloid):

not quite. let's break it down into two steps: this is the left side: x this is the right side: 3x + 6 now SUBTRACT 3x from the left side, what do you get?

Vocaloid (vocaloid):

remember how we added + 2 before? it's the same thing, except you SUBTRACT 3x instead of ADDING 2

OpenStudy (anonymous):

would x= become -3x?

Vocaloid (vocaloid):

Not quite, what is x minus 3x?

OpenStudy (anonymous):

3x

Vocaloid (vocaloid):

no, what is x - 3x?

Vocaloid (vocaloid):

easier question, what is 1 - 3?

OpenStudy (anonymous):

-2

Vocaloid (vocaloid):

right! so what do you think x - 3x is?

OpenStudy (anonymous):

are you saying that it is -2x?

Vocaloid (vocaloid):

YES! we're done with the left side, let's move to the right side: 3x + 6 now subtract 3x and tell me what you get!

OpenStudy (anonymous):

6x? 3x+6=9x 9x-6x=(3x)?

Vocaloid (vocaloid):

I'm not sure where you're getting 9x "subtract 3x" means minus 3 x 3x + 6 - 3x = ?

OpenStudy (anonymous):

6x

Vocaloid (vocaloid):

no, please try again

Vocaloid (vocaloid):

easier question: what is a + b - a?

OpenStudy (anonymous):

b

Vocaloid (vocaloid):

Right! so what is 3x + 6 - 3x?

OpenStudy (anonymous):

2?

Vocaloid (vocaloid):

no, please try again... easier question: what is 3x - 3x?

OpenStudy (anonymous):

0x

Vocaloid (vocaloid):

right, and 0x is the same as 0, right? so what is 3x + 6 - 3x?

Vocaloid (vocaloid):